# Photonic Simulation of Localization Phenomena Using Boson Sampling

### v1

@@ II.1 Theoretical Overview of Boson Sampling
|$\ket{\psi}=\ket{n_{1}^{(I)},n_{2}^{(I)},...,n_{M}^{(I)}}=\prod_{j=1}^{M}\frac{(\hat{a_{j}})^{n_{j}^{(I)}}}{[n_{j}^{(I)}!]^{1/2}}\ket{\textbf{0}},$ (1)
@@
where $F$ indexes all the possible configurations of the final state, $n_{i}^{(F)}$ represents the number of photons in the $i^{th}$ mode for configuration $S$, and $\gamma_{F}$ is the amplitude associated with the configuration $F$. The probability $P_{F}$ of obtaining a particular configuration $F$ will be given by $P_{F}=|\gamma_{F}|^{2}$. The probability amplitudes of $N$ particle boson sampling are given by permanents of $N\times N$ sub-matrices of the $M\times M$ interferometer unitary. These sub-matrices are constructed with rows chosen according to the input state and columns chosen according to the output state, with repeated copies of the corresponding row/column in case of coincidence events. An example of a boson sampling circuit and the relation between probabilities and permanents (Eq. (4)) is displayed in Fig. (1). For a specific configuration, this looks like
$P_{F}=\lvert\gamma_{F}\rvert^{2}=\frac{\lvert\mathrm{Per}[U^{F,I}]\rvert^{2}}{\prod_{i=1}^{M}n_{i}^{(I)}!\prod_{i=1}^{M}n_{i}^{(F)}!}.$ (4)
@@ IV.1 Results: Phase Transition in AAH Model
Figure 6: Dynamics of AAH model for varying values of quasi-periodic disorder strength $\lambda$. (a) Dynamics of second-order participation entropy $S_{2}^{PE}(t)$, normalized by the natural logarithm of the system size ($M=7$). (b) Dynamics of Inverse Participation Ratio $I_{2}$(t). The data for each plot was computed for 30 values of $\lambda$, with the evolution upto $t=40$ discretized into 100 timesteps, and averaged over 50 realizations of $\delta$.
@@
$I_{2}(t)=\sum_{i}^{M}|P_{i}(t)|^{2}.$ (26)
IPR helps to quantify the spread of the wavefunction over all the energy eigenstates. This is especially useful for wavefunctions in superpositions over multiple states. For normalized wavefunctions in an extended state, $I_{2}\sim 1/\sqrt{\mathcal{N}}$, whereas for a strongly localized state, $I_{2}\sim 1$, with intermediate behaviour in the critical phase[99, 100, 65, 55, 101].
In Fig. 6 (a), we illustrate the dynamics of the second order participation entropy $S_{2}^{PE}$, upto late times. We can see that after a fast initial relaxation, the participation entropy in all cases shows an oscillatory behaviour around a particular value. For smaller values of the quasiperiodic disorder strength ($\lambda<<2$), we can see that $S_{2}^{PE}/\log{\mathcal{N}}$ ) tends to oscillate close to 1, denoting the delocalized phase. As $\lambda$ increases, transport is suppressed and $S_{2}^{PE}$ at late times is bounded by a constant, oscillating at a much lower value. This is a signature of the system being in the localized phase.
@@ Acknowledgments
We are grateful to the High Performance Computing (HPC) facility at IISER Bhopal, where large-scale calculations in this project were run. A.K is grateful to I-HUB Quantum Technology Foundation, IISER Pune for her Chanakya Post-Graduate Fellowship [Ref. I-HUB/PGF/B-3/2023-24/07]. V.T is grateful to DST-INSPIRE for her PhD fellowship. A.S acknowledges financial support from SERB via the grant (File Number: CRG/2019/003447), and from DST via the DST-INSPIRE Faculty Award [DST/INSPIRE/04/2014/002461]. This research made use of Strawberry Fields version 0.23.0, an open-source software framework for photonic quantum computing.
@@ Appendix A On the effect of sample count
It should be noted that in an experimental scenario, we will have to perform multiple “shots” of the same circuit each with a photon-counting measurement following the evolution. Performing an average over the samples thus obtained will be necessary to approximate the Fock space probabilities. A higher sample count therefore leads to greater accuracy of the result. Let us consider an example to illustrate this effect. Consider the time evolution of a photon in the critical phase of the Aubry-André-Harper model. Fig. 9 depicts a comparison between five instances of the same circuit, with the first four columns showing results for an increasing number of shots, while the fifth column showing results via the all_fock_probs method. It can be seen in Figure 9 (a) that the trend gradually nears the expected behaviour as the number of shots is increased.
@@
We can see in Fig. (9) (b), that as the number of shots increases, $\Delta$ becomes lesser, which implies a higher accuracy of the simulated result. One may also use other metrics such as state-fidelity to study the accuracy in such a case.

### v2

@@ II.1 Theoretical Overview of Boson Sampling
|$\ket{\psi}=\ket{n_{1}^{(I)},n_{2}^{(I)},...,n_{M}^{(I)}}=\prod_{j=1}^{M}\frac{(\hat{a}_{j}^{\dagger})^{n_{j}^{(I)}}}{[n_{j}^{(I)}!]^{1/2}}\ket{\textbf{0}},$ (1)
@@
where $F$ indexes all the possible configurations of the final state, $n_{i}^{(F)}$ represents the number of photons in the $i^{th}$ mode for configuration $S$, and $\gamma_{F}$ is the amplitude associated with the configuration $F$. The probability $P_{F}$ of obtaining a particular configuration $F$ will be given by $P_{F}=|\gamma_{F}|^{2}$. The probability amplitudes of $N$ particle boson sampling are given by permanents of $N\times N$ sub-matrices of the $M\times M$ interferometer unitary. These sub-matrices are constructed with rows chosen according to the input state and columns chosen according to the output state, with repeated copies of the corresponding row/column in case of coincidence events. If $U^{(F,I)}$ denotes the sub-matrix for a specific configuration, the probability of obtaining that configuration is given by
$P_{F}=\lvert\gamma_{F}\rvert^{2}=\frac{\lvert\mathrm{Per}[U^{(F,I)}]\rvert^{2}}{\prod_{i=1}^{M}n_{i}^{(I)}!\prod_{i=1}^{M}n_{i}^{(F)}!}.$ (4)
An example of a boson sampling circuit and the relation between probabilities and permanents (Eq. (4)) is displayed in Fig. (1).
@@ IV.1 Results: Phase Transition in AAH Model
Figure 6: Dynamics of AAH model for varying values of quasi-periodic disorder strength $\lambda$. (a) Dynamics of second-order participation entropy $S_{2}^{PE}(t)$, normalized by the natural logarithm of the system size ($M=7$). (b) Dynamics of Inverse Participation Ratio $I_{2}$(t). The data for each plot was computed for evolution upto $t=40$ discretized into 100 timesteps, and averaged over 50 realizations of $\delta$.
@@
$I_{2}(t)=\sum_{i}^{M}P_{i}(t)^{2}.$ (26)
IPR helps to quantify the spread of the wavefunction over all the energy eigenstates. This is especially useful for wavefunctions in superpositions over multiple states. For normalized wavefunctions in an extended state, $I_{2}\sim 1/\sqrt{\mathcal{N}}$, whereas for a strongly localized state, $I_{2}\sim 1$, with intermediate behaviour in the critical phase[99, 100, 65, 101, 93, 55].
In Fig. 6 (a), we illustrate the dynamics of the second order participation entropy $S_{2}^{PE}$, upto late times. We can see that after a fast initial relaxation, the participation entropy in all cases shows an oscillatory behaviour around a particular value. For smaller values of the quasiperiodic disorder strength ($\lambda<2$), we can see that $S_{2}^{PE}/\log{\mathcal{N}}$ ) tends to oscillate close to 1, denoting the delocalized phase. As $\lambda$ increases, transport is suppressed and $S_{2}^{PE}$ at late times does continues to oscillate around a much lower value and does not exceed some constant upper limit – this constant value depends on the magnitude of $\lambda$, and is independent of the size of the configuration space. This is a signature of the system being in the localized phase.
@@ Acknowledgments
We are grateful to the High Performance Computing (HPC) facility at IISER Bhopal, where large-scale calculations in this project were run. A.K acknowledges funding support for Chanakya - PG fellowship from the National Mission on Interdisciplinary Cyber Physical Systems, of the Department of Science and Technology, Govt. of India through the I-HUB Quantum Technology Foundation. V.T is grateful to DST-INSPIRE for her PhD fellowship. A.S acknowledges financial support from SERB via the grant (File Number: CRG/2019/003447), and from DST via the DST-INSPIRE Faculty Award [DST/INSPIRE/04/2014/002461]. This research made use of Strawberry Fields version 0.23.0, an open-source software framework for photonic quantum computing.
@@ Appendix A On the effect of sample count
It should be noted that in an experimental scenario, we will have to perform multiple “shots” of the same circuit each with a photon-counting measurement following the evolution. Performing an average over the samples thus obtained will be necessary to approximate the Fock space probabilities. A higher sample count therefore leads to greater accuracy of the result. Let us consider an example to illustrate this effect. Consider the time evolution of a photon in the critical phase of the Aubry-André-Harper model. Fig. (9) depicts a comparison between five instances of the same circuit, with the first four columns showing results for an increasing number of shots, while the fifth column showing results via the all_fock_probs() method. It can be seen in Figure 9 (a) that the trend gradually nears the expected behaviour as the number of shots is increased.
@@
We can see in Fig. 9 (b), that as the number of shots increases, $\Delta$ becomes lesser, which implies a higher accuracy of the simulated result. This decrease in the error margin is also apparent in the probability density of the state. We show in Fig. (10) a comparison of this across multiple shot values. One may also use other metrics such as state-fidelity to study the accuracy in such a case.
[…]
Figure 10: Comparison of probability density of the wave-function at five representative timestamps, for varying shot counts. We consider the exact result as the one obtained from the all_fock_probs()— method. The system is an AAH model, with state at $t=0$ initialized to |$\ket{0001000}$, and system parameters $J=1,\delta=0,M=7,\lambda=2$. The probability estimation at 10 shots per timestep is poor, however we can see that as the number of shots increases, the error margin $\Delta$ reduces significantly.